OpenStudy (anonymous):

How to do factorization? i.e. 6x^2 - 8x + 4 = 0 ? And to do the 'complete the square' on it as well?

OpenStudy (kainui):

So to do this one, we will divide both sides of the equation by 6 to get the first term to just be x^2. You'll want to do this most of the time, since it'll make your life easier. See, what we're going to do is try to make it fit into a simple squared formula like this: (x+m)^2 So if you ever have trouble figuring it out, just write that, and then multiply the two to get: x^2+2mx+m^2=(x+m)^2 Alright, now if you look at what you have, you'll have something multiplied by x. Make that equal to 2m. Now you'll be able to solve for m and square it, then add it to both sides of the formula. So for example if you have: x^2+14x-6=0 you can see that 14 is where the 2m is on the other thing. So, 14=2m m=7 m^2=49 So what you can do to make your life a little easier is add 6 to both sides so it looks like this, and we'll use that m^2 in a second. x^2+14x-6=0 add 6 to both sides x^2+14x=6 now if you add 49 (m^2) to both sides you get: x^2+14x+49=55 We chose m^2 very carefully so that now we can go from: x^2+2m+m^2 to (x+m)^2 since they're equal, right? x^2+14x+49=55 Now we do the step (x+7)^2=55 So take the square root of both sides \[x+7=\pm \sqrt{55}\]\[x=-7\pm \sqrt{55}\] So those are our two possible answers. The reason for the + and - sign on the square root is because if you take any number, like -2 or 2 and square them, they both equal 4, so there are two possible choices. That's how you complete the square basically.